YOU DRAW 2 CARDS FROM A STANDARD DECK WITHOUT REPLACEMENT. DETERMINE THE PROBABILITY THAT:
A. BOTH ARE DIAMONDS
B. NEITHER IS A DIAMOND
C. BOTH ARE OF THE SAME SUIT
D. BOTH ARE OF THE SAME SUIT, GIVEN THAT THE FIRST DRAWN IN A CLUB
E. THEY ARE NOT OF THE SAME SUIT
F. THEY ARE NOT OF THE SAME SUIT, GIVEN THAT THE FIRST DRAWN IN A CLUB.

Answer by stanbon(75887)

YOU DRAW 2 CARDS FROM A STANDARD DECK WITHOUT REPLACEMENT. DETERMINE THE PROBABILITY THAT:
A. P(BOTH ARE DIAMONDS) = 13C2/52C2 = 78/1326 = 0.0588...
------------------------------------------------------------------
B. EITHER IS A DIAMOND = 1 - 0.0588 = 0.9412
------------------------------------------------------------------
C. BOTH ARE OF THE SAME SUIT = 4*0.0588 = 0.2353
------------------------------------------------------------------
D. BOTH ARE OF THE SAME SUIT, GIVEN THAT THE FIRST DRAWN IN A CLUB
P(C|C| = (13/52) * (12/51) = 0.0588...
------------------------------------------------------------------
E. THEY ARE NOT OF THE SAME SUIT = 1 - 0.0588 = 0.9412..
----------------------------------------------------------
F. THEY ARE NOT OF THE SAME SUIT, GIVEN THAT THE FIRST DRAWN IS A CLUB.
P(not club|Club) = 12/51*13/52 = 0.0588
================================================
Cheers,
Stan H.
====================